Angles X and Y are complementary. If mX = a – 12 and mY = a + 8, find the measure of mX. Please help me, I am doing pre-algebra and my teacher is very strict. This is yesterday's "problem of the day" and almost everybody got it.
Is the answer 70? Please respond with haste.
Thanks

Answer by mananth(16946)

(a-12)+(a+8)=90
2a-4=90
2a= 94
a=47
m X = a-12==> 47-12= 35
mY= a+8==> 47+8=55